OpenStudy (anonymous):

I need help with this question! t - 2(3 - 2t) = 2t + 9

OpenStudy (anonymous):

(t -2)(3 - 2t) = 2t + 9 2t^2 + 7t - 6 = 2t + 9

OpenStudy (anonymous):

thats as far as I have got

OpenStudy (anonymous):

@cj49 @dan815 @mathmath333 @Jumperman

OpenStudy (princeharryyy):

2t^2+5t - 15=0 solve using t = (-b+sqrt(D))/2a or t =(-b-sqrt(D))/2a

OpenStudy (anonymous):

????

OpenStudy (anonymous):

its imaginary number i guess

OpenStudy (anonymous):

I think maybe I set it up wrong....

OpenStudy (cj49):

t-6+4t=2t+9 3t-2t=9+6 t=15

OpenStudy (anonymous):

ok so first get your t's to one side so subtract 1t from the left side. then do distributive property on the -2(3-2t) part so thats -6+4t=1t+9 then subtract 1t from 4t so thats -6+3t+9 then add your 6 to your 9 so thats 15 then you have 3t=15 and 15/3 =5

OpenStudy (anonymous):

t=5

OpenStudy (princeharryyy):

@cj49 what she she did is (t -2)(3 - 2t) = 2t + 9

OpenStudy (princeharryyy):

so u can't put it that way @cj49

OpenStudy (anonymous):

ok I just checked my answer book and 5 is the right answer... @stevietomat0 was right

OpenStudy (cj49):

bt i guess this is the question ryt "t - 2(3 - 2t) = 2t + 9" @princeharryyy

OpenStudy (anonymous):

i was right!!!

OpenStudy (anonymous):

I realize I should have only multiplied 2 by (3 - 2t) ....

OpenStudy (princeharryyy):

she later corrected that @cj49 if she meant wat u solved then its obviously ohk :))

OpenStudy (anonymous):

cool should put me as a best answer, im new to all of this xD

OpenStudy (anonymous):

thanks everyone for the help! :) @stevietomat0 @cj49 @princeharryyy

OpenStudy (anonymous):

just did :) @stevietomat0

OpenStudy (cj49):

lol i messed up i din subtracted the t first like stevietomat0 did

OpenStudy (anonymous):

lol its ok :) thank u for the help !! @cj49

OpenStudy (cj49):

lol np c:..